A body travel 100 m in the first 2 seconds and 104 m in the next 4 seconds .How far will it move in the next 4 seconds of the acceleration is uniform

Grade:11

1 Answers

Sudeep
7 years ago
From eqn. S=ut+1/2at^2For t=2,S=100 , solving this we get a relation between u &a as u+a=50.Acc. is uniform and it moves 104m in next 4 secAgain using the same eqn. Taking S=204(from the starting point) in time T=6sec .We get u+3a=36 . Solving these eqns we get u= 57 & a= -7.So the displacement covered in next 4 sec is S(at 10 sec from starting point)-S(at 6 sec) = -84 m.[But if you meant how much "distance they travel" u also need to consider the stopping point (I.e at 8.1428 sec for this case and at 232m for this prob. You get distance covered=140m)